OpenStudy (anonymous):

(4,8) m=2 in point slope form

OpenStudy (anonymous):

y = mx + b

sam (.sam.):

\[y-y_1=m(x-x_1) \\ \\ Substitute (4,8),m=2, \\ \\ y-8=2(x-4) \\ \\ y=2x-8+8 \\ \\ y=2x\]